OpenStudy (anonymous):

A coal car on a train weighs 30 tons plus 1 ton per cubic yard of coal x that it carries. The total weight of a coal car is: f(x) = x + 30. How will the graph of this function change if the coal car weight is changed to 26 tons? A.)The line will shift vertically up by 4 tons. B.) The line will shift vertically up 26 tons. C.) The line will shift vertically down by 4 tons. D.) The line will shift vertically down by 26 tons.

OpenStudy (anonymous):

Okay Cassie. So 30 is the added weigh in the equation. Or the y-int where it starts. So if you compare f(x)=x+30 to f(x)=x+26 What changes?

OpenStudy (anonymous):

I goes down 4

OpenStudy (anonymous):

Yes. Vertically. becuse it's the y -intercept that's changing.

OpenStudy (anonymous):

So the answer is C. Good job.

OpenStudy (anonymous):

Thank you